Prove that 2^(n+1) > (n + 2) · sin(n) for all positive integers n


1
Expert's answer
2021-10-07T11:37:00-0400

Let "P(n)" be the proposition

Basis Step. "P(1)" is true because

"2^{1+1}=4>3=1+2."

Inductive Step.

Assume that "P(k)" holds for an arbitrary positive integer "k." That is, we assume that


"2^{k+1}> k + 2"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"2^{(k+1)+1}> (k+1) + 2"

"2^{(k+1)+1}=2(2^{k+1})"

"=2^{k+1}+2^{k+1}>k+2+k+2"

"=k+3+k+1>k+3"

This shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers "n." That is, we have proved that "2^{n+1}> n + 2" for all positive integers "n."


"\sin(n)\leq1, n\in\R"

Then for all positive integers "n"


"(n+2)\cdot\sin(n)\leq n+2, n\in\R"

We have proved that "2^{n+1}> n + 2" for all positive integers "n."

Therefore we have proved that

"2^{n+1}>(n+2)\cdot\sin(n)"

for all positive integers "n."